OpenStudy (anonymous):

what is the value of 6*8-4[7+(8-3)]?

OpenStudy (anonymous):

I keep getting 0 im I correct

OpenStudy (alexandervonhumboldt2):

ok let me do

OpenStudy (alexandervonhumboldt2):

6*8-4[7+(8-3)]=48-4*(7+5)=48-4*12=0 correct

OpenStudy (anonymous):

Thank You AlexanndervonHumboldt2